# bibliography.bib

### v1

@@ Abstract
Weight sharing, equivariance, and local filters, as in convolutional neural networks, are believed to contribute to the sample efficiency of neural networks. However, it is not clear how each one of these design choices contribute to the generalization error. Through the lens of statistical learning theory, we aim to provide an insight into this question by characterizing the relative impact of each choice on the sample complexity. We obtain lower and upper sample complexity bounds for a class of single hidden layer networks. It is shown that the gain of equivariance is directly manifested in the bound, while getting a similar increase for weight sharing depends on the sharing mechanism. Among our results, we obtain a completely dimension-free bound for equivariant networks for a class of pooling operations. We show that the bound depends merely on the norm of filters, which is tighter than using the spectral norm of the respective matrix. We also characterize the trade-off in sample complexity between the parametrization of filters in spatial and frequency domains, particularly when spatial filters are localized as in vanilla convolutional neural networks.
@@ 1 Introduction
In recent years, equivariant neural networks have gained particular interest within the machine learning community thanks to their inherent ability to preserve certain transformations in the input data, thereby providing a form of inductive bias that aligns with many real-world problems. Indeed, equivariant networks have shown remarkable performance in various applications, ranging from computer vision to molecular chemistry, where data often exhibit specific forms of symmetry. Equivariant networks are closely related to their more traditional counterparts, Convolutional Neural Networks. CNNs are a particular class of neural networks that achieve equivariance to translation. The convolution operation in CNNs ensures that the response to a particular feature is the same, regardless of its spatial location in the input data. However, the concept of equivariance extends beyond just translation, accommodating a broader spectrum of transformations such as rotations, reflections, and scaling, among others. Group Convolutional Neural Networkss [cohen_group_2016] are the typical example of equivariant neural networks. The inductive bias introduced by equivariance, in the form of symmetry preservation, offers an intuitive connection to their generalization capabilities. By encoding prior knowledge about the structure of the data, equivariant networks can efficiently exploit the inherent symmetries, reducing sample complexity and improving generalization performance. This ability to generalize from a limited set of examples is crucial to their success. Their mathematical foundation is grounded in group representation theory, which provides a powerful framework to understand and design neural networks that are equivariant or invariant to the action of a group of transformations.
Besides the inductive bias of data symmetry, the generalization benefits of CNNs and GCNNs are attributed additionally to the weight-sharing, implemented by the convolution operation, and the locality, implemented by the smaller filter size. In this paper, we study the impact of equivariance, weight-sharing, and locality on generalization within the framework of statistical learning theory. Our focus will be on neural networks with one hidden layer. Similar to works like [vardi_sample_2022, magen_initialization-dependent_2023], we believe that this study provides a first step toward a better understanding of deeper networks. Getting dimension-free generalization error bounds constitutes an important line of research in the literature. Following this line of work, we provide various dimension-free and norm-based bounds for one hidden layer networks. See Appendix J.1 for discussions on the desiderata of learning theory.
@@
We consider a class of one hidden layer networks where the first layer is a multi-channel equivariant layer followed by point-wise non-linearity, a pooling layer, and the final linear layer. We assume that the $\ell_{2}$-norm of the parameters of each layer is bounded. For architectures based on group convolution with point-wise non-linearity, we provide generalization bounds for various pooling operations that are completely dimension-free. We obtain a similar norm-based for general equivariant networks. We also provide a lower bound on Rademacher complexity that shows the tightness of our bound. We show that no gain is observed if the analysis is conducted for the networks parameterized in the frequency domain. When the equivariant layer is replaced by a layer with, not necessarily equivariant, weight-sharing, we provide as well a dimension-free bound. By studying the bound, it can be seen that some particular weight-sharing schemes, although not all of them, can provide similar generalization guarantees. Next, we provide another bound for networks with local filters and show that the locality can bring additional gain on top of equivariance. We will then show that there is a trade-off between locality in the spatial domain and frequency domain, important for band-limited inputs. The trade-off is characterized by the uncertainty principle. Finally, we provide the numerical verification of our bounds. The generalization bounds are all obtained using Rademacher complexity analysis. We relegate all proofs to the appendix.
@@ 2 Preliminaries
Figure 1: A group convolutional network with $c_{0}$ input channel, $c_{1}$ intermediate channels, and a last layer with pooling and linear aggregation ${\bm{u}}$.
We first introduce key concepts from group and representation theories necessary to present our main results. In the rest of this work, we will generally assume a compact group $G$. Note that this includes any finite group. For the most part, we will consider Abelian groups, i.e., commutative groups (such as planar rotations or periodic translations).
@@
Note also that the group $G$ naturally acts on a signal $x:G\to\mathbb{R}^{c}$ as $g:x\mapsto g.x$, with $[g.x](h)=x(g^{-1}h)$. Then, one can show that the group convolution operator above is equivariant with respect to this action on its input and output, i.e. $w\circledast_{G}g.x=g.(w\circledast_{G}x)$. Additionally, it is well known, e.g. [generaltheory, Kondor2018-GENERAL], that group convolutions with learnable filters in this form parameterize the most general linear equivariant maps between feature spaces of signals over a compact group $G$. Finally, GCNN typically uses an activation function applied point-wise on the convolution output. In our 1-layer architecture, a first convolution layer is followed by a pointwise activation layer and, then, a per-channel - average or max - pooling layer, which produces $c_{1}$ invariant features. These final features are mixed with the final linear layer ${\bm{u}}$. See Figure 1 for more details. The final network is then given as:
@@ 3 Related Works
Previous works attempted to improve machine learning models by leveraging prior knowledge about the symmetries and the geometry of a problem [bronstein2021geometric]. A variety of design strategies have been explored in the literature to achieve group equivariance, for example via equivariant MLPs [Shawe-Taylor_1993, Shawe-Taylor_1989, finzi2021practical], group convolution [cohen_group_2016, Kondor2018-GENERAL, bekkers2018roto], Lie group convolution [bekkers2020bspline, finzi_generalizing_2020], steerable convolution [cohen_steerable_2016, generaltheory, Worrall2017-HNET, Weiler2018-STEERABLE, 3d_steerableCNNs, TensorFieldNets, Weiler2019, NEURIPS2020_15231a7c, brandstetter2021geometric, cesa2022a] and, very recently, by using geometric algebra [ruhe2023clifford, brehmer2023geometric], to only mention a few. Other previous approaches include [mallat_group_2012, Dieleman2016-CYC, defferrard_deepsphere_2019]. Some of these ideas have also been used to generalize convolution beyond groups to more generic manifolds via the framework of Gauge equivariance [cohen2019gauge, weiler_coordinate_2021]. While equivariance is generally considered a powerful inductive bias leading to improved data efficiency, this large selection of equivariant designs raises the following question: what impact different architectural choices have on performance and to what extent do they aid generalization?
@@
Some previous works tried to answer some aspects of this question. [taylowThresholdLearning] is of the first works studying the relation between invariance and generalization. For example, [sokolic_generalization_2017] extend the robustness-based generalization bounds found in [sokolic_robust_2017] by assuming the set of transformations of interest change the inputs drastically, thereby proving that the generalization error of a $G$-invariant classifier scales $1/\sqrt{|G|}$, where $|G|$ is the cardinality of the finite equivariance group $G$. [bietti2019stability] study the stability of some particular compact group equivariant models and also describe an associated Rademacher complexity and generalization bound. [lyle2020benefits] investigate the effect of invariance under the lenses of the PAC-Bayesian framework but do not provide an explicit bound. Later, [elesedy_provably_2021] use VC-dimension analysis and derive more concrete bounds. [elesedy_group_2022] studies compact-group equivariance in PAC learning framework and equates learning with equivariant hypotheses and learning on a reduced data space of orbit representatives to obtain a sample complexity bound. [sannai_improved_2021] propose a similar idea, proving that equivariant models work in a reduced space, the Quotient Feature Space (QFS), whose volume directly affect the generalization error. While the result relaxes the robustness assumption in [sokolic_generalization_2017], the final bound is suboptimal with respect to the sample size. PAC learnability under transformation invariance is also studied in [shao2022theory]. [zhu_understanding_2021] characterize the generalization benefit of invariant models using an argument based on the covering number induced by a set of transformations. More recently, [behboodi_pac-bayesian_2022] leverages a representation-theoretic construction of equivariant networks and provides a norm-based PAC-Bayesian generalization bound inspired by [neyshabur_pac-bayesian_2018]. Finally, [petrache2023approximation] considers a more general setting, allowing for approximate, partial, and misspecified equivariance and studying how the relation between data and model equivariance error impacts generalization.
@@
Many works in the literature have previously investigated the generalization properties of deep learning methods. [zhang_understanding_2017] first noted that a model trained on random labels can achieve small training error while producing arbitrarily large generalization errors. This result raised a new challenge in the field since popular uniform complexity measures such as VC-dimensions are inconsistent with this finding. This inspired many recent works which tried to explain generalization in terms of other quantities, such as margin or norms of the learnable weights; a few nonexhaustive examples are [wei_improved_2019, sokolic_robust_2017, neyshabur_pac-bayesian_2018, arora_stronger_2018, bartlett_spectrally-normalized_2017, golowich_size-independent_2018, dziugaite_entropy-sgd_2018, long_size-free_2019, vardi_sample_2022, ledent_norm-based_2021, valle-perez_generalization_2020]. The PAC Bayesian framework is a particularly popular method and has been applied to neural networks in many previous works, e.g. see [neyshabur_pac-bayesian_2018, biggs_non-vacuous_2022, dziugaite_computing_2017, dziugaite_data-dependent_2018, dziugaite_entropy-sgd_2018, dziugaite_search_2020, lotfi2022pac]. [jiang_fantastic_2020] perform a thorough experimental comparison of many complexity measures and identifies some failure cases; see also [nagarajan_uniform_2019, koehler_uniform_2021, negrea_defense_2021] for further discussion on uniform complexity measures and their limitations. The norm-based bounds can still be tight and informative for shallow networks considered in this work.
@@
As one of the most popular deep learning architectures, convolutional neural networks (CNNs) have received particular attention in the literature. Generalization studies on CNNs are especially relevant for our work since we focus significantly on equivariance to finite and abelian groups, which can often be realized via periodic convolution. [pitas_limitations_2019, long_size-free_2019, vardi_sample_2022, ledent_norm-based_2021] previously studied these architectures. In particular, [graf_measuring_2022] recently established two covering-numbers-based bounds for convolution models.
@@ 4 Sample Complexity Bounds For Equivariant Networks
We consider the group convolutional networks as defined in 3. We assume that the input to the network is bounded as $\left\|{\bm{x}}\right\|\leq b_{x}$ [^1] For all the results in the paper, we can use a data-dependent bound on the input. Namely, it suffices to assume that $\max_{i\in[m]}\left\|{\bm{x}}_{i}\right\|\leq b_{x}$.. Consider the following hypothesis space:
@@
Since the pooling operation is a permutation invariant operation, we can use Theorem 7 in [zaheer_deep_2017] to show that one can always find two functions $\phi:\mathbb{R}^{|G|+1}\to\mathbb{R}$ and $\rho:\mathbb{R}\to\mathbb{R}^{|G|+1}$ such that $P\circ{\bm{z}}=\phi\left(\frac{1}{|G|}\sum_{i=1}^{|G|}\rho(z_{i})\right)$. We can provide the following generalization bound for a subset of such representations, namely for positive homogeneous $\phi:\mathbb{R}\to\mathbb{R}$ and $\rho:\mathbb{R}\to\mathbb{R}$.
@@
Consider the hypothesis space ${\mathcal{H}}$ defined in 4. If $P(\cdot)$ is the pooling operation represented as $P\circ{\bm{z}}=\phi(\frac{1}{|G|}\mathbf{1}^{\top}\rho({\bm{z}}))$, where the two functions $\rho(\cdot)$, $\phi(\cdot)$ and the activation function $\sigma(\cdot)$ are all $1$-Lipschitz positive homogeneous activation function, then with probability at least $1-\delta$ and for all $h\in{\mathcal{H}}$, we have:
@@
The proof leverages Rademacher complexity analysis and is presented in Appendix C.4. Note that the assumption of positive homogeneity is only needed to utilize the peeling technique of [golowich_size-independent_2018]. We expect that similar techniques from [vardi_sample_2022] can be used to extend the result to Lipschitz activation functions.
@@
Consider now the special case of average pooling operation $P({\bm{x}})=\frac{1}{|G|}\mathbf{1}^{\top}{\bm{x}}$. This is a linear layer; therefore, it can be combined with the last layer ${\bm{u}}$ to yield a standard two-layer neural network. We can utilize the results from [vardi_sample_2022, golowich_size-independent_2018]. The combination of the layer ${\bm{u}}$ and average pooling is a linear layer with the norm $M_{1}/\sqrt{|G|}$, and the matrix ${\bm{W}}$, a circulant matrix, has the Frobenius norm $\sqrt{|G|}M_{2}$. The product of these norms would be bounded by $M_{1}M_{2}$. Being a special case of the model in [vardi_sample_2022], we can use their results off-the-shelf to get an upper bound on the sample complexity that depends only on $M_{1}M_{2}$ for Lipschitz-activation functions (Theorem 2 of [vardi_sample_2022]). We provide an independent proof for positive homogeneous activation functions in Appendix C.2 with a clean form. Note that the authors in [vardi_sample_2022] provide a result for average pooling that contains the term $O_{\phi}$, the maximal number of patches that any single input coordinate appears in. This term is equal to $|G|$ in our case, which is canceled out by the average pooling term. Their bound depends on the spectral norm of the underlying circulant matrix, which can be bigger than the norm of the filter provided here. But, their proof can be reworked with the norm of convolutional filter instead, in which case, their result will be the special case of ours.
@@
We cannot rely on the previous results for Max pooling operation, since the peeling argument would not work. Theorem 7 and 8 of [vardi_sample_2022] provide a bound for max pooling, however their network does not contain the linear aggregation ${\bm{u}}$ after max-pooling, and their bound contains dimension dependencies such as $\log(|G|c_{1})$ or $\log(m)$. In Appendix C.3, we provide a bound for max pooling that depends merely on $\log(|G|)$ but contains a data dependent term that can have poor scaling.
@@
The authors in [behboodi_pac-bayesian_2022] improved the PAC-Bayesian generalization bound by conducting their analysis using the representation in the frequency domain. In our Rademacher analysis, such a shift would not bring any additional gain, and we basically recover the same bound. We provide the details in the supplementary materials.

### v2

@@ Abstract
Weight sharing, equivariance, and local filters, as in convolutional neural networks, are believed to contribute to the sample efficiency of neural networks. However, it is not clear how each one of these design choices contributes to the generalization error. Through the lens of statistical learning theory, we aim to provide insight into this question by characterizing the relative impact of each choice on the sample complexity. We obtain lower and upper sample complexity bounds for a class of single hidden layer networks. For a large class of activation functions, the bounds depend merely on the norm of filters and are dimension-independent. We also provide bounds for max-pooling and an extension to multi-layer networks, both with mild dimension dependence. We provide a few takeaways from the theoretical results. It can be shown that depending on the weight-sharing mechanism, the non-equivariant weight-sharing can yield a similar generalization bound as the equivariant one. We show that locality has generalization benefits, however the uncertainty principle implies a trade-off between locality and expressivity. We conduct extensive experiments and highlight some consistent trends for these models.
@@ 1 Introduction
In recent years, equivariant neural networks have gained particular interest within the machine learning community thanks to their inherent ability to preserve certain transformations in the input data, thereby providing a form of inductive bias that aligns with many real-world problems. Indeed, equivariant networks have shown remarkable performance in various applications, ranging from computer vision to molecular chemistry, where data often exhibit specific forms of symmetry. Equivariant networks are closely related to their more traditional counterparts, Convolutional Neural Networks. CNNs are a particular class of neural networks that achieve equivariance to translation. The convolution operation in CNNs ensures that the response to a particular feature is the same, regardless of its spatial location in the input data. However, equivariance extends beyond just translation, accommodating a broader spectrum of transformations such as rotations, reflections, and scaling. Group Convolutional Neural Networkss [cohen_group_2016] are the typical example of equivariant neural networks. The inductive bias introduced by equivariance, in the form of symmetry preservation, offers an intuitive connection to their generalization capabilities. By encoding prior knowledge about the structure of the data, equivariant networks can efficiently exploit the inherent symmetries, reducing sample complexity and improving generalization performance. This ability to generalize from a limited set of examples is crucial to their success. Their mathematical foundation is grounded in group representation theory, which provides a powerful framework to understand and design neural networks that are equivariant or invariant to the action of a group of transformations.
Besides the inductive bias of data symmetry, the generalization benefits of CNNs and GCNNs are additionally attributed to the weight-sharing implemented by the convolution operation and the locality implemented by the smaller filter size. In this paper, we study the impact of equivariance, weight-sharing, and locality on generalization within the framework of statistical learning theory. Our focus will be on neural networks with one hidden layer. Similar to works like [vardi_sample_2022, magen_initialization-dependent_2023], we believe that this study provides a first step toward a better understanding of deeper networks. Getting dimension-free generalization error bounds constitutes an important line of research in the literature. Following this line of work, we provide various dimension-free and norm-based bounds for one hidden layer networks. See Appendix J.1 for discussions on the desiderata of learning theory.
@@
We consider a class of one hidden layer networks where the first layer is a multi-channel equivariant layer followed by point-wise non-linearity, a pooling layer, and the final linear layer. We assume that the $\ell_{2}$-norm of the parameters of each layer is bounded. For architectures based on group convolution with point-wise non-linearity, we provide generalization bounds for various pooling operations that are entirely dimension-free. We obtain a similar norm-based bound for general equivariant networks. We also provide a lower bound on Rademacher complexity that shows the tightness of our bound. We extend the results to max-pooling, combining various covering number arguments for Rademacher complexity analysis. The bound is only dimension-dependent on the number of hidden layer channels and logarithmic-dependent on the group size; otherwise, it is independent of other dimensions. We also extend the result to multi-layer networks and discuss its limitations. We show that no gain is observed if the analysis is conducted for the networks parameterized in the frequency domain. When a layer replaces the equivariant layer with, not necessarily equivariant, weight-sharing, we also provide a dimension-free bound. By studying the bound, it can be seen that some particular weight-sharing schemes, although not all, can provide similar generalization guarantees. Next, we give another bound for networks with local filters and show that the locality can bring additional gain on top of equivariance. We will then show a trade-off between locality in the spatial and frequency domains, which is important for band-limited inputs. The uncertainty principle characterizes the trade-off. Finally, we provide the numerical verification of our bounds. The generalization bounds are all obtained using Rademacher complexity analysis. We relegate all proofs to the appendix.
@@ 2 Preliminaries
Figure 1: Visualization of the network architectures with equivariance, locality, and weight sharing. On the right, we also summarize how each choice impacts the generalization error in our theory.
Convolution: a simple example.
We start with a simple example of familiar convolutional neural networks, indulging some of the subtleties needed for our later discussions. For an RGB image input, each convolution filter will slide over the input image and act on its receptive field by simple multiplication. If we translate the pixels of the input image, ignoring the corner pixels, the output of the previous convolution operation would just get translated as well. The act of translation is an example of a group action, and the fact that this act shifts the convolution output is an example of equivariance property [^1] We would like to emphasize again that the example is not perfect. First, the convolution in CNN is a cross-correlation. Second, the conventional CNNs are only approximately translation equivariant because of the finite input size and the corner pixels.. In what follows, we work with a generalized notion of action captured by group theory, which studies objects like translation and rotation that can combine and have an inverse element.
We introduce key concepts from group and representation theories necessary to present our main results in Appendix A. In the rest of this work, we will generally assume a compact group $G$. Note that this includes any finite group. For the most part, we will consider Abelian groups, i.e., commutative groups (such as planar rotations or periodic translations).
@@
Note also that the group $G$ naturally acts on a signal $x:G\to\mathbb{R}^{c}$ as $g:x\mapsto g.x$, with $[g.x](h)=x(g^{-1}h)$. Then, one can show that the group convolution operator above is equivariant with respect to this action on its input and output, i.e., $w\circledast_{G}g.x=g.(w\circledast_{G}x)$. Additionally, it is well known, e.g,. [generaltheory, Kondor2018-GENERAL], that group convolutions with learnable filters in this form parameterize the most general linear equivariant maps between feature spaces of signals over a compact group $G$. Finally, GCNN typically uses an activation function applied point-wise on the convolution output. In our 1-layer architecture, a first convolution layer is followed by a pointwise activation layer and, then, a per-channel - average or max - pooling layer, which produces $c_{1}$ invariant features. These final features are mixed with the final linear layer ${\bm{u}}$. See Figure 1 for more details. The final network is then given as:
@@ 3 Related Works
Previous works attempted to improve machine learning models by leveraging prior knowledge about the symmetries and the geometry of a problem [bronstein2021geometric]. A variety of design strategies have been explored in the literature to achieve group equivariance, for example via equivariant MLPs [Shawe-Taylor_1993, Shawe-Taylor_1989, finzi2021practical], group convolution [cohen_group_2016, Kondor2018-GENERAL, bekkers2018roto], Lie group convolution [bekkers2020bspline, finzi_generalizing_2020], steerable convolution [cohen_steerable_2016, generaltheory, Worrall2017-HNET, Weiler2018-STEERABLE, 3d_steerableCNNs, TensorFieldNets, Weiler2019, NEURIPS2020_15231a7c, brandstetter2021geometric, cesa2022a] and, very recently, by using geometric algebra [ruhe2023clifford, brehmer2023geometric], to only mention a few. Other previous approaches include [mallat_group_2012, Dieleman2016-CYC, defferrard_deepsphere_2019]. Some of these ideas have also been used to generalize convolution beyond groups to more generic manifolds via the framework of Gauge equivariance [cohen2019gauge, weiler_coordinate_2021]. While equivariance is generally considered a powerful inductive bias that improves data efficiency, this large selection of equivariant designs raises the following question: What impact do different architectural choices have on performance, and to what extent do they aid generalization?
@@
Some previous works tried to answer some aspects of this question. [taylowThresholdLearning] is one of the first works studying the relation between invariance and generalization. For example, [sokolic_generalization_2017] extend the robustness-based generalization bounds found in [sokolic_robust_2017] by assuming the set of transformations of interest change the inputs drastically, thereby proving that the generalization error of a $G$-invariant classifier scales $1/\sqrt{|G|}$, where $|G|$ is the cardinality of the finite equivariance group $G$. [bietti2019stability] study the stability of some particular compact group equivariant models and also describe an associated Rademacher complexity and generalization bound. [lyle2020benefits] investigate the effect of invariance under the lenses of the PAC-Bayesian framework but do not provide an explicit bound. Later, [elesedy_provably_2021] use VC-dimension analysis and derive more concrete bounds. [elesedy_group_2022] studies compact-group equivariance in PAC learning framework and equates learning with equivariant hypotheses and learning on a reduced data space of orbit representatives to obtain a sample complexity bound. [sannai_improved_2021] propose a similar idea, proving that equivariant models work in a reduced space, the Quotient Feature Space (QFS), whose volume directly affects the generalization error. While the result relaxes the robustness assumption in [sokolic_generalization_2017], the final bound is suboptimal with respect to the sample size. PAC learnability under transformation invariance is also studied in [shao2022theory]. [zhu_understanding_2021] characterize the generalization benefit of invariant models using an argument based on the covering number induced by a set of transformations. More recently, [behboodi_pac-bayesian_2022] leverages a representation-theoretic construction of equivariant networks and provides a norm-based PAC-Bayesian generalization bound inspired by [neyshabur_pac-bayesian_2018]. Finally, [petrache2023approximation] considers a more general setting, allowing for approximate, partial, and misspecified equivariance and studying how the relation between data and model equivariance error impacts generalization.
@@
Many works in the literature have previously investigated the generalization properties of deep learning methods. [zhang_understanding_2017] first noted that a model trained on random labels can achieve small training errors while producing arbitrarily large generalization errors. This result raised a new challenge in the field since popular uniform complexity measures such as VC dimensions are inconsistent with this finding. This inspired many recent works which tried to explain generalization in terms of other quantities, such as margin or norms of the learnable weights; a few non-exhaustive examples are [wei_improved_2019, sokolic_robust_2017, neyshabur_pac-bayesian_2018, arora_stronger_2018, bartlett_spectrally-normalized_2017, golowich_size-independent_2018, dziugaite_entropy-sgd_2018, long_size-free_2019, vardi_sample_2022, ledent_norm-based_2021, valle-perez_generalization_2020]. The PAC Bayesian framework is a particularly popular method. It has been applied to neural networks in many previous works, e.g. see [neyshabur_pac-bayesian_2018, biggs_non-vacuous_2022, dziugaite_computing_2017, dziugaite_data-dependent_2018, dziugaite_entropy-sgd_2018, dziugaite_search_2020, lotfi2022pac]. [jiang_fantastic_2020] perform a thorough experimental comparison of many complexity measures and identifies some failure cases; see also [nagarajan_uniform_2019, koehler_uniform_2021, negrea_defense_2021] for further discussion on uniform complexity measures and their limitations. The norm-based bounds can still be tight and informative for shallow networks considered in this work.
@@
As one of the most popular deep learning architectures, convolutional neural networks (CNNs) have received particular attention in the literature. Generalization studies on CNNs are especially relevant for our work since we focus significantly on equivariance to finite and abelian groups, which can often be realized via periodic convolution. [pitas_limitations_2019, long_size-free_2019, vardi_sample_2022, ledent_norm-based_2021] previously studied these architectures. The authors in [vardi_sample_2022] represented a convolutional layer as a linear layer applied on local patches and used Rademacher complexity to get the bound. In [long_size-free_2019], the bound is derived using Vapnik-Chervonenkis analysis [Vapnik2015-yp, Gine2001-jj] and depends on the number of parameters but independent of the number of pixels in the input. In [graf_measuring_2022], the authors use two covering-numbers-based bounds for Rademacher complexity analysis of convolutional models. Please see Appendix J for further discussions.
@@ 4 Sample Complexity Bounds For Equivariant Networks
We consider the group convolutional networks as defined in eq. 3. We assume that the input to the network is bounded as $\left\|{\bm{x}}\right\|\leq b_{x}$ [^2] For all the results in the paper, we can use a data-dependent bound on the input. Namely, it suffices to assume that $\max_{i\in[m]}\left\|{\bm{x}}_{i}\right\|\leq b_{x}$.. Consider the following hypothesis space:
@@
Since the pooling operation is a permutation invariant operation, we can use Theorem 7 in [zaheer_deep_2017] to show that one can always find two functions $\phi:\mathbb{R}^{|G|+1}\to\mathbb{R}$ and $\rho:\mathbb{R}\to\mathbb{R}^{|G|+1}$ such that $P\circ{\bm{z}}=\phi\left(\frac{1}{|G|}\sum_{i=1}^{|G|}\rho(z_{i})\right)$. We can provide the following generalization bound for a subset of such representations, namely for positively homogeneous $\phi:\mathbb{R}\to\mathbb{R}$ and $\rho:\mathbb{R}\to\mathbb{R}$.
@@
Consider the hypothesis space ${\mathcal{H}}$ defined in eq. 4. If $P(\cdot)$ is the pooling operation represented as $P\circ{\bm{z}}=\phi(\frac{1}{|G|}\mathbf{1}^{\top}\rho({\bm{z}}))$, where the two functions $\rho(\cdot)$, $\phi(\cdot)$ and the activation function $\sigma(\cdot)$ are all $1$-Lipschitz positively homogeneous activation function, then with probability at least $1-\delta$ and for all $h\in{\mathcal{H}}$, we have:
@@
The proof leverages Rademacher complexity analysis and is presented in Appendix C.5. Note that the assumption of positively homogeneity is only needed to utilize the peeling technique of [golowich_size-independent_2018]. We expect that similar techniques from [vardi_sample_2022] can be used to extend the result to Lipschitz activation functions.
@@
Consider now the special case of average pooling operation $P({\bm{x}})=\frac{1}{|G|}\mathbf{1}^{\top}{\bm{x}}$. This is a linear layer; therefore, it can be combined with the last layer ${\bm{u}}$ to yield a standard two-layer neural network. We can utilize the results from [vardi_sample_2022, golowich_size-independent_2018]. The combination of the layer ${\bm{u}}$ and average pooling is a linear layer with the norm $M_{1}/\sqrt{|G|}$, and the matrix ${\bm{W}}$, a circulant matrix, has the Frobenius norm $\sqrt{|G|}M_{2}$. The product of these norms would be bounded by $M_{1}M_{2}$. Being a special case of the model in [vardi_sample_2022], we can use their results off-the-shelf to get an upper bound on the sample complexity that depends only on $M_{1}M_{2}$ for Lipschitz-activation functions (Theorem 2 of [vardi_sample_2022]). We provide an independent proof for positively homogeneous activation functions in Appendix C.2 with a clean form. Note that the authors in [vardi_sample_2022] provide a result for average pooling that contains the term $O_{\phi}$, the maximal number of patches that any single input coordinate appears. In our case, this term equals $|G|$, which is canceled out by the average pooling term. Their bound depends on the spectral norm of the underlying circulant matrix, which can be bigger than the norm of the filter provided here. But, their proof can be reworked with the norm of convolutional filter instead, in which case, their result will be our special case.
@@
We cannot rely on the previous results for the max pooling operation since the peeling argument would not work. Theorem 7 and 8 of [vardi_sample_2022] provide a bound for max pooling, however their network does not contain the linear aggregation ${\bm{u}}$ after max-pooling, and their bound contains dimension dependencies such as $\log(|G|c_{1})$ or $\log(m)$. We provide various bounds for max pooling in Appendix C.3 and C.4. The proof technique is different from the above results. Again, the bounds have different dimension dependencies. We believe these dependencies are proof artifacts. Removing them would be an interesting direction for future work.
@@
The authors in [behboodi_pac-bayesian_2022] improved the PAC-Bayesian generalization bound by conducting their analysis using the representation in the frequency domain. In our Rademacher analysis, such a shift would not bring any additional gain, and we recover the same bound. We provide the details in the supplementary materials.
4.1 Bounds for Multi-Layer Equivariant Network
In this section, we study a simple extension to multi-layer group equivariant networks. Consider the following network with $L$ hidden layer:
${h_{{\bm{u}},\{{\bm{w}}^{(l)},l\in[L]\}}({\bm{x}}):={\bm{u}}^{\top}P\circ{\sigma({\bm{W}}^{(L)}\sigma({\bm{W}}^{(L-1)}\dots\sigma({\bm{W}}^{(1)}){\bm{x}}\dots)}}$ (5)
where $P(\cdot)$ is the average pooling operation, $\sigma(\cdot)$ is the ReLU function. Each linear layer ${\bm{W}}^{l}$ is the same as the mapping in eq. 2 but with $c_{l-1}$ and $c_{l}$ as the input and output channels. The new hypothesis space is given by:
${{\mathcal{H}}^{(L)}:=\left\{h_{{\bm{u}},\{{\bm{w}}^{(l)},l\in[L]\}}:\left\|{\bm{u}}\right\|\leq M_{1},\left\|{\bm{w}}_{i}\right\|\leq M_{i+1},i\in[L]\right\}.}$ (6)
We have the following theorem.
Theorem 4.3.
Consider the hypothesis space ${\mathcal{H}}^{(L)}$ defined in eq. 6. With probability at least $1-\delta$ and for all $h\in{\mathcal{H}}$, we have:
${\mathcal{L}(h)\leq\hat{\mathcal{L}}(h)+2\frac{b_{x}|G|^{\frac{L-1}{2}}M_{1}M_{2}\dots M_{L+1}}{\sqrt{m}}+4\sqrt{\frac{2\log(4/\delta)}{m}}.}$
The proof can be found in Appendix C.6. We first comment on the dimension dependency of the bound. Although our bound has no dependence on the number of input and hidden layer channels, it still has dependence on $|G|^{(L-1)/2}$. It is not entirely dimension-free. Using some other techniques, for example, see Section 3 in [golowich_size-independent_2018], it can be improved to $(L-1)\log|G|$ for $L>1$. However, dimension dependence is only one concern of the above result. It is worth pointing out that the generalization bounds for deeper networks suffer from many shortcomings and generally fail to correlate well with the empirical generalization error. For example, various norm bounds were considered in [jiang_fantastic_2020], and the correlation with the generalization error was explored. Generally, the norm-based bounds performed poorly, while sharpness-aware bounds showed promises. Besides, in [nagarajan_uniform_2019], the norm-based bounds were shown to increase drastically with the training set size, leading to looser bounds with increasing training set size. Therefore, we think these norm-based bounds have limitations in deep network generalization analysis.